OpenStudy (anonymous):

Prove identity: tan^2 (x) - sin^2 (x) =tan^2(x)sin^2(x)

OpenStudy (dumbcow):

use following identities: tan^2 = sin^2/cos^2 sin^2 +cos^2 =1 sin^2/cos^2 - sin^2 = (sin^2 - sin^2cos^2)/cos^2 =sin^2(1-cos^2)/cos^2 =sin^2/cos^2 * (1-cos^2) = tan^2*sin^2